# Bayesian Low-Rank LeArning (Bella): A Practical Approach to Bayesian Deep Learning

### v3

@@ Abstract
Computational complexity of Bayesian learning is impeding its adoption in practical, large-scale tasks, despite demonstrations of significant merits such as improved robustness and resilience to unseen or out-of-distribution inputs over their non-Bayesian counterparts. Although Deep ensemble methods (Seligmann et al. 2024; Lakshminarayanan, Pritzel, and Blundell 2017) have proven to be highly effective for Bayesian deep learning, their practical application is hindered by substantial computational cost. In this study, we introduce an innovative framework to mitigate the computational burden of ensemble Bayesian deep learning. We explore a more feasible alternative, inspired by the recent success of low-rank adapters, we introduce Bayesian Low-Rank LeArning (Bella). We show, i) Bella achieves a dramatic reduction in the number of trainable parameters required to approximate a Bayesian posterior; and ii) it not only maintains, but in some instances, surpasses the performance–in accuracy and out-of-distribution generalisation–of conventional Bayesian learning methods and non-Bayesian baselines. Our extensive empirical evaluation in large-scale tasks such as ImageNet, CAMELYON17, DomainNet, VQA with CLIP, LLaVA demonstrate the effectiveness and versatility of Bella in building highly scalable and practical Bayesian deep models for real-world applications.
@@ 1 Introduction
Figure 1: Train and Test Error (%) landscapes of CAMELYON17. Training landscape demonstrates the learned 3 particle approximation of the modes of the posterior from a pre-trained model $\boldsymbol{\theta}$—modification of parameters in the constrained region ($\Delta\boldsymbol{\theta}$) leads to approaching posterior modes in inference. Interestingly, the test landscape shows a shift in particle performance, reflecting the intrinsic distribution shift. Here, we observe a key benefit of our Bella approximation compared to a point estimate—while a single parameter particle (e.g. $\boldsymbol{\theta}+\Delta\boldsymbol{\theta}_{1}$) do not generalize well, Bayesian prediction with Equation 1, effectively an average over multiple parameter settings, leads to better performance.
Bayesian deep learning (Neal 2012) provides mechanisms, for building predictive models more robust to adversarial attacks, resilient to unseen or out-of-distribution data and a theoretical framework for estimating model uncertainty (Liu et al. 2019; Ye and Zhu 2018; Wilson et al. 2022; Izmailov et al. 2020; Chen and Ghattas 2020). In particular, quantifying uncertainty enables more reliable decision-making and facilitates identifying potential model vulnerabilities in uncertain regions of the input space. Consequently, embracing Bayesian deep learning (BDL) represents a significant stride towards building more reliable and trustworthy AI systems for various real-world applications (e.g., autonomous driving, medical image analysis, etc.). Unfortunately, their practical use is encumbered by their computational complexity.
Unlike traditional alternatives with point estimates—a single set of model parameters mapping inputs to outputs—Bayesian models learn the distribution of model parameters to offer a distribution over possible predictions. Consider a neural network $f(\mathbf{x},\boldsymbol{\theta})$ with input $\mathbf{x}$ parameterized by $\boldsymbol{\theta}$ and its corresponding prior distribution $p(\boldsymbol{\theta})$. The likelihood $p(\mathcal{D}|\boldsymbol{\theta})$ is determined by $f$ (Bishop 2006) and then, Bayesian inference seeks to use Bayes’ theorem to derive a posterior distribution given by $p\left(\boldsymbol{\theta}|\mathcal{D}\right)={p\left(\mathcal{D}|\boldsymbol{\theta}\right)p\left(\boldsymbol{\theta}\right)}/{p(\mathcal{D})}$ to compute the predictive distribution:
@@
Despite evidence to support the adoption of Bayesian learning, their widespread use is hindered by several challenges; a primary issue is the intractability of the posterior distribution $p(\boldsymbol{\theta}|\mathcal{D})$ in practical learning tasks. The exact solution for the posterior, even for networks of moderate size, is impractical. Because of deep neural network’s complexity and the high-dimensional integral of the resulting denominator. This necessitates the use of approximations.
Our work expands upon recent research demonstrating the effectiveness of ensemble methods and Stein Variational Gradient Descent (SVGD) (Liu and Wang 2016) approaches in BDL (e.g., (Seligmann et al. 2024; Lakshminarayanan, Pritzel, and Blundell 2017; Abbasnejad et al. 2020a; Doan et al. 2022; Izmailov et al. 2020; Chen and Ghattas 2020)) to provide a practical method for ensemble Bayesian deep learning for large-scale tasks. To achieve a practical and efficient approach, we consider spawning parameter particles (models) by adding linear interpolation to a pre-trained model’s parameters to an ensemble. To further enhance the efficiency of these interpolations, inspired by the success of low-rank adapters (LORAs)(Hu et al. 2022), we propose using low-rank parameters in the interpolation process, called Bayesian Low-Rank Learning (Bella).
To validate this, we compare Bella to the full-parameter alternative in terms of (1) accuracy, (2) out-of-distribution generalization, (3) adversarial robustness, and (4) alignment of uncertainty estimation with human confidence. In summary, our findings indicate that this approach achieves comparable performance to full SVGD or ensemble methods at a fraction of the computational cost, with the exception of adversarial robustness. For example, Figure 1 illustrates the application of our proposed approach of Bayesian inference in an application with distribution shifts (CAMELYON17 (Koh et al. 2021) benchmark). The particles (i.e., parameter samples from the posterior) obtained using Bella prove effective in improving generalization, improving the common single particle fine-tuning and comparable with full ensemble-based baselines at a fraction of the cost. Our key contributions are:
@@
  • We demonstrate Bella to perform on par or better than baselines—ensembles or current SVGD implementations. Bella consistently outperforms the non-BNN counterparts on 8 datasets; image classification (ImageNet, CIFAR 10/100), Out-of-distribution (DomainNet, Camleyon17, CIFAR-10-C), and VQA as well as adversarial robustness. Notably, it sets a new state-of-the-art for the CAMELYON17 (see Table 7 for more details) according to the leaderboard (Challenge 2024).
  • Bella employed with the multi-modal model LlaVa (Liu et al. 2023b), leads to improved performance and uncertainty estimation highly correlated with human confidence.
@@ 2 Related Work
3 Background on Stein Variational Gradient Descent (SVGD)
@@ 5.6 Ablation Studies
5.7 Generalization to a Visual Question Answer
@@
Notably, our method is efficient, particularly in contrast to the resource-intesive baseline Bayesian models tailored for this task (e.g. (Abbasnejad et al. 2019, 2020b, 2020a)). The computational requirements for using these baselines render the application of full SVGD or ensemble alternatives impractical, thereby highlighting the impact on practical applications by harnessing the effectiveness of Bella.
@@ 6 Conclusion
In this paper, we introduce an innovative approach for creating an efficient Bayesian Neural Network (BNN) approximation using only a base pre-trained model. Our approach, called Bella, demonstrates remarkable compatibility with full-rank BNN approximations like SVGD or Ensemble methods, outperforming single-network solutions across a range of tasks, from pure classification to out-of-distribution (OOD) generalization and uncertainty quantification. Our research paves the way for effective BNN implementation, facilitating the development of reliable and robust machine-learning models capable of scaling to large networks and datasets. Bella reinforces the effectiveness of employing such a simple, efficient and effective training method with diverse representations over the conventional single fine-tuning techniques prevalent in the field.
@@ A Practical Approach to Bayesian Neural Networks
    Additional comparison studies. Including SoTA results on (Challenge 2024) (Appendix E)
@@ Appendix B Generalization of CIFAR-10 Models on OOD dataset (STL-10)
Experimental Set-up. We choose eight labels in STL-10 test set, same labels observed in CIFAR-10 dataset to set a OOD dataset. This includes 6,400 images acting as the OOD dataset for CIFAR-10 pre-trained models.
@@ Appendix E Additional Comparison Studies
Table 7: Choosing different layers for fine-tuning with Bella, with $r=4,n=5$ on CAMELYON17 achieves state-of-the-art performance (Challenge 2024) (see highlighted row).

### v4

@@ Abstract
Bayesian learning has demonstrated merits in robustness and resilience to unseen and out-of-distribution inputs, yet its adoption in large-scale tasks remains limited. In particular, deep ensemble methods (Seligmann et al. 2024; Lakshminarayanan, Pritzel, and Blundell 2017) have proven to be highly effective, but they suffer from a high computational cost. This paper introduces an alternative framework that mitigates the computational burden of ensemble Bayesian deep learning. The approach is inspired by the recent success of low-rank adapters and is named Bayesian Low-Rank LeArning (Bella). We show that (1) Bella achieves a dramatic reduction in the number of trainable parameters required to approximate a Bayesian posterior; and (2) it not only maintains, but in some instances surpasses the performance—in accuracy and out-of-distribution generalisation—of conventional Bayesian and non-Bayesian methods. An extensive empirical evaluation on large-scale tasks (ImageNet, CAMELYON17, DomainNet, VQA) with CLIP and LLaVA demonstrates the effectiveness and versatility of Bella in building scalable and Bayesian deep models for real-world applications.
@@ 1 Introduction
Bayesian deep learning (Neal 2012) provides mechanisms for building predictive models that are more robust to adversarial attacks, resilient to unseen and out-of-distribution inputs, and a principled approach for estimating model uncertainty (Liu et al. 2019; Ye and Zhu 2018; Wilson et al. 2022; Izmailov et al. 2020; Chen and Ghattas 2020). In particular, quantifying uncertainty enables more reliable decision-making and facilitates identifying potential model vulnerabilities in uncertain regions of the input space. Therefore, Bayesian deep learning (BDL) has a high potential in helping build more reliable and trustworthy AI systems for various real-world applications (e.g., autonomous driving, medical image analysis, etc.). Unfortunately, BDL methods suffer from a high computational complexity.
Bayesian models contrast with traditional point estimates, i.e. a single set of parameters representing a single mapping from inputs to outputs. A Bayesian model, instead, learns a distribution of model parameters that captures a distribution over possible predictions. Let us consider a neural network $f(\mathbf{x},\boldsymbol{\theta})$ with input $\mathbf{x}$ parameterized by $\boldsymbol{\theta}$ and its corresponding prior distribution $p(\boldsymbol{\theta})$. The likelihood $p(\mathcal{D}|\boldsymbol{\theta})$ is determined by $f$ (Bishop 2006) and Bayesian inference uses Bayes’ theorem to derive a posterior distribution $p\left(\boldsymbol{\theta}|\mathcal{D}\right)={p\left(\mathcal{D}|\boldsymbol{\theta}\right)p\left(\boldsymbol{\theta}\right)}/{p(\mathcal{D})}$ and then compute the predictive distribution:
@@
Figure 1: Training (left) and test (right) error landscapes on the CAMELYON17 dataset (%). The left figure shows a learned 3-particle approximation of the modes of the posterior of a pre-trained model $\boldsymbol{\theta}$. The test landscape is slightly different from the training one, highlighting the challenge of generalization. While a point estimate does not generalize well (a single parameter particle such as $\boldsymbol{\theta}+\Delta\boldsymbol{\theta}_{1}$) Bella shows a clear benefit. The Bayesian prediction (Equation 1) effectively averages over multiple parameter settings and leads to lower test error.
Despite evidence for the benefits of Bayesian learning, its adoption is hindered by several challenges, most importantly the intractability of the posterior distribution $p(\boldsymbol{\theta}|\mathcal{D})$ in most practical tasks. The exact solution is impractical to compute even for networks of moderate size, hence the necessity of using approximations.
This paper expands upon recent research demonstrating the effectiveness of ensemble methods and Stein Variational Gradient Descent (SVGD) (Liu and Wang 2016) approaches in BDL (Seligmann et al. 2024; Lakshminarayanan, Pritzel, and Blundell 2017; Abbasnejad et al. 2020a; Doan et al. 2022; Izmailov et al. 2020; Chen and Ghattas 2020). These methods aim to provide a practical method for ensemble Bayesian deep learning for large-scale tasks. With the goal of achieving a practical and efficient approach, we consider spawning parameter particles (models) by adding linear interpolation to a pre-trained model’s parameters to an ensemble. To further enhance the efficiency of these interpolations, we take inspiration from low-rank adaptation (LoRAs)(Hu et al. 2022) and propose to use low-rank parameters in the interpolation process. We therefore name the approach Bayesian Low-Rank Learning (Bella).
We validated Bella by comparing it to the full-parameter alternative in terms of (1) accuracy, (2) out-of-distribution generalization, (3) adversarial robustness, and (4) alignment of uncertainty estimation with human confidence. In summary, our findings indicate that our approach achieves comparable performance to full SVGD or ensemble methods at a fraction of the computational cost, with the exception of adversarial robustness. For example, Figure 1 illustrates the application of our proposed approach of Bayesian inference in an application with distribution shifts (CAMELYON17 (Koh et al. 2021) benchmark). The particles (i.e., parameter samples from the posterior) obtained using Bella prove effective in improving generalization, improving the common single particle fine-tuning and comparable with full ensemble-based baselines at a fraction of the cost. Our key contributions are:
@@
  • We demonstrate Bella to perform on par or better than baselines—ensembles or current SVGD implementations. Bella consistently outperforms the non-BNN counterparts on 8 datasets; image classification (ImageNet, CIFAR 10/100), Out-of-distribution (DomainNet, Camleyon17, CIFAR-10-C), and VQA as well as adversarial robustness. Notably, it sets a new state-of-the-art on CAMELYON17 (see Table 7 and the Camelyon leaderboard (2024)).
  • Bella employed with the multi-modal model LlaVa (Liu et al. 2023b) leads to improved performance and uncertainty estimation highly correlated with human confidence.
@@ 2 Related Work
3 Background on
Stein Variational Gradient Descent (SVGD)
@@ 5.6 Ablation Studies
5.7 Visual Question Answering
@@
Notably, our method is efficient, particularly in contrast to the resource-intesive baseline Bayesian models tailored for this task, e.g. Abbasnejad et al. (2019, 2020b, 2020a). The computational requirements for using these baselines render the application of full SVGD or ensemble alternatives impractical, thereby highlighting the impact on practical applications by harnessing the effectiveness of Bella.
@@ 6 Conclusion
In this paper, we introduce an innovative approach for creating an efficient Bayesian Neural Network (BNN) approximation using only a base pre-trained model. Our approach, called Bella, demonstrates remarkable compatibility with full-rank BNN approximations like SVGD or Ensemble methods, outperforming single-network solutions across a range of tasks, from pure classification to out-of-distribution (OOD) generalization and uncertainty quantification. Our research paves the way for effective and efficient BNNs, facilitating the development of reliable and robust machine-learning models. Bella reinforces the effectiveness of employing such a simple, efficient and effective training method with diverse representations over the conventional practice of single fine-tuning.
@@ A Practical Approach to Bayesian Neural Networks
    Additional comparison studies. Including SoTA results on (Camelyon leaderboard 2024) (Appendix E)
@@ Appendix B Generalization of CIFAR-10 Models on OOD dataset (STL-10)
Experimental Setup. We choose eight labels in STL-10 test set, same labels observed in CIFAR-10 dataset to set a OOD dataset. This includes 6,400 images acting as the OOD dataset for CIFAR-10 pre-trained models.
@@ Appendix E Additional Comparison Studies
Table 7: Choosing different layers for fine-tuning with Bella, with $r=4,n=5$ on CAMELYON17 achieves state-of-the-art performance (Camelyon leaderboard 2024) (see highlighted row).